### v2

# Standardizing Generative Face Video Compression using Supplemental Enhancement Information

Abstract
This paper proposes a Generative Face Video Compression (GFVC) approach using Supplemental Enhancement Information (SEI), where a series of compact spatial and temporal representations of a face video signal (i.e., 2D/3D key-points, facial semantics and compact features) can be coded using SEI message and inserted into the coded video bitstream. At the time of writing, the proposed GFVC approach using SEI messages has been adopted into the official working draft of Versatile Supplemental Enhancement Information (VSEI) standard by the Joint Video Experts Team (JVET) of ISO/IEC JTC 1/SC 29 and ITU-T SG16, which will be standardized as a new version for “ITU-T H.274 ISO/IEC 23002-7”. To the best of the authors’ knowledge, the JVET work on the proposed SEI-based GFVC approach is the first standardization activity for generative video compression. The proposed SEI approach has not only advanced the reconstruction quality of early-day Model-Based Coding (MBC) via the state-of-the-art generative technique, but also established a new SEI definition for future GFVC applications and deployment. Experimental results illustrate that the proposed SEI-based GFVC approach can achieve remarkable rate-distortion performance compared with the latest Versatile Video Coding (VVC) standard, whilst also potentially enabling a wide variety of functionalities including user-specified animation/filtering and metaverse-related applications.
Index Terms:
Generative AI, face video coding, supplemental enhancement information, VVC, JVETI Introduction
Recently, generative AI technology has greatly accelerated the progress of Generative Face Video Compression (GFVC) [1, 2, 3, 4], showing promising performance potentials and application diversities over traditional codecs [5, 6]. Different from early-day Model-Based Coding (MBC) techniques [7, 8], GFVC exploits the strong inference capabilities of deep generative models to improve the face reconstruction quality using very compact spatial and temporal representations (i.e., 2D/3D key-points, facial semantics, compact feature and others). Inspired by promising rate-distortion performance and the potential to enable more diverse applications, the Joint Video Experts Team (JVET) of ISO/IEC JTC 1/SC 29 and ITU-T SG16 has established an Ad hoc Group (AhG) for comprehensive GFVC investigations, including coordinated experimentation, software implementation, interoperability and so on. In the past year, the JVET GFVC AhG has established GFVC test conditions [9] and software tools [10, 11], and the Generative Face Video (GFV) Supplemental Enhancement Information (SEI) message has been included into the official working draft [12] of Versatile Supplemental Enhancement Information (VSEI) standard and will be finally standardized in “ITU-T H.274 ISO/IEC 23002-7”.
From the perspective of JVET standardization activities, the proposed approach provides a reasonable solution to a standardize GFVC with SEI messages, bridging deep generative models with face video compression tasks using standardized bitstream format. In particular, standardizing GFVC with an SEI message allows various GFVC feature formats to be carried and combines these features seamlessly with the bitstreams from a traditional hybrid video codec without introducing normative changes to the encoder/decoder of the hybrid video codec. As a result, the proposed SEI-based GFVC standardization approach can effectively utilize advanced generative AI techniques to provide performance gains and functionalities over traditional codecs. The main contributions of this paper can be summarized as follows,
The potential to standardize generative video compression with SEI messages has been explored in JVET. Taking face video as an example, a novel SEI-based standardization approach is proposed to specify compact facial representations using a unified high-level syntax structure that can be inserted into the coded video bitstream of a hybrid video codec.
The proposed SEI message for generative face video can well support different facial representations, such as 2D/3D landmarks, 2D/3D keypoints, compact feature and facial semantics. Moreover, this GFV SEI message can also provide flexibility for different functionalities such as ultra-low bitrate communication, user-specified animation/filtering, and metaverse-related applications.
The proposed GFV SEI message has been implemented on the top of the VVC reference software and extended with common GFVC software tools, which were both endorsed by JVET as GFVC AhG activities. As such, the entire SEI-based GFVC approach can be executed in a streamlined manner. The simulation results demonstrate the effectiveness of the proposed SEI-based GFVC approach, which can achieve promising compression performance and diverse functionalities over the VVC standard.
II Related Work
II-A Generative Face Video Compression
In terms of ultra-low bitrate visual compression, parameterized face video compression can be tracked to early MBC techniques [7, 8] in the 1990s, which provided promising compression capability via statistical priors and image synthesis. However, limited by the model learning and generalization abilities at that time, the reconstruction quality was not sufficient for the early-day MBC techniques to be further developed and applied in the following years. Recently, inspired by the strong inference capabilities of deep generative models like Generative Adversarial Network (GAN) [13] and Diffusion Model (DM) [14], the MBC concept has been revived and further extended into a novel generative compression paradigm. Taking face video as an example, deep generative models can facilitate the characterization of a high-dimensional face signal into compact features, and use these features to perform accurate motion estimation for high-quality face video reconstruction.
In particular, keypoint-based GFVC frameworks [2] can realize ultra-low bit-rate face video communication by compressing a series of learned 2D keypoints along with their local affine transformations. A wide variety of facial representations like 2D facial landmarks [15], 3D keypoints [3], compact temporal features [4] and interactive facial semantics [16] have been proposed to support more robust and versatile GFVC algorithms. However, these algorithms still simply bridge generative animation models with compression tasks, and lack more mature consideration such as how to achieve commercial deployment and enable real applications through standardization. As for the requirements of GFVC standardization, it should focus on specifying the bitstream format for a variety of GFVC feature representations in a unified form and ensuring the interoperability/robustness of GFVC systems. Moreover, standardized GFVC format can realize ultra-low bitrate face video communication and improve the quality of experience of face video applications like video conferencing/chat. It also potentially enables a wide variety of functionalities including user-specified animation/filtering and metaverse-related applications.
II-B Supplemental Enhancement Information
As additional data specified into the bitstream, SEI messages can convey extra information and support various post-processing technology. In the past decades, JVET has successfully developed many high-level syntax structures [17, 18] to enable versatile scenarios and new functionalities like viewport-adaptive 360° immersive media.
From the year of 2023, JVET has started to investigate the standardization of GFVC, covering topics from how to carry various GFVC features using standardized syntax to how to insert GFV SEI messages into the VVC coded bitstream. Specifically, a common GFV SEI message syntax [19, 20] was proposed to support the carriage of different GFVC feature formats within a VVC coded bitstream, whilst the interfaces to GFVC translator, generator, and enhancer neural networks [21, 22, 23] were also specified. Interoperability among different GFVC feature formats [24], fusion/enhancement modules for quality improvement [25, 26, 27] and model complexity reduction for practical deployment [28] were further investigated to enable GFVC applications towards better performance. GFVC software tools [10, 11] have been developed to support various GFVC algorithms, and the GFV SEI message and its enhancement extension have also been included into the official working draft [12] of VSEI standard and will be finally standardized in “ITU-T H.274 ISO/IEC 23002-7”. Herein, this paper summarizes how to enable different GFVC representations with GFV SEI message. The paper also describes a technical implementation to link GFVC software tools with GFV SEI message implementation for comprehensive performance evaluations.
III SEI-based Generative Compression Approach
In this section, we provide the GFV SEI message definition and its workflow in pseudo-code format regarding how to encapsulate different GFVC feature formats with a unified GFV SEI message syntax. We also discuss how to support interoperability among different GFVC feature formats and improve reconstruction robustness via fusion/enhancement techniques. Finally, we describe the detailed encoding and decoding processes for the proposed SEI-based GFVC approach.
| GFVC Methods | Facial Representation | Parameter Format |
| VSBNet [15] | 2D Landmarks | Coordinate |
| FOMM [29] | 2D Keypoints+Affine Transformation Matrices | Coordinate & Matrix |
| DAC [2] | 2D Keypoints | Coordinate |
| FV2V [3] | 3D Keypoints+Head Rotation/Translation Matrices | Coordinate & Matrix |
| Mob M-SPADE [30] | Segmentation Map | Matrix |
| SNRVC [31] | Facial Semantics | Matrix |
| CFTE [4] | Compact Feature Matrices | Matrix |
| CTTR [32] | Compact Feature Matrices | Matrix |
| Bi-Net [33] | 2D Keypoints | Coordinate |
| RDAC [34] | 2D Keypoints + Residual Map | Coordinate & Matrix |
| IFVC [16] | Facial Semantics (Mouth/Eye/Head Rotation/Translation Matrices) | Matrix |
III-A GFV SEI Message Design
A face signal exhibits strong statistical regularities, which can be efficiently characterized with 2D landmarks, 2D keypoints, 3D keypoints, compact features, facial semantics or other formats. Such compact facial description strategies can lead to reduced coding bit-rate and improved coding efficiency for GFVC algorithms. However, diverse facial representations also make it difficult to standardize GFVC. Herein, we summarize a few typical GFVC algorithms and attempt to discover the characteristics of their compact representations. As shown in Table I, these compact facial representations can be classified into two categories: coordinate-related parameters and matrix-related parameters. In addition, all GFVC algorithms proposed in the literature use a hybrid codec to code the base pictures. These observations confirm that there is great potential to standardize GFVC. Intuitively, a SEI message provides a reasonable standardized way to accommodate different GFVC representations into a unified format and convey these compact representations together with VVC-coded base pictures. As such, SEI-based GFVC approach has recently been a subject active study in JVET.
Algorithm 1 provides a pseudocode representing the syntax structure of GFV SEI message according to the GFV SEI message text that has been included in the working draft of VSEI version 4 [12]. In particular, the GFV SEI message indicates facial parameters and specifies a facial parameter translator network, denoted as TranslatorNN, and a face picture generator neural network, denoted as GenerativeNN. Besides, the GFV SEI message provides specific flags/factors to determine parameter type, whether prediction is used, and codeword length. It should be emphasized that the GFV SEI message generator in Algorithm 1 is just a brief summary of the GFV SEI message included in the working draft of VSEI version 4 [12]. Readers are referred to the detailed syntax and semantics of GFV SEI message draft text in [12] for further information, including 2D/3D parameters, number of coordinates, matrix dimensions and flags for extra functionalities, as well as the usage interface for TranslatorNN and GenerativeNN that use these GFV parameters for signal reconstruction.
III-B GFVC System Optimization
The proposed GFV SEI message semantics for the VSEI standard provides a specification of syntax and semantics sufficient for usage, but does not establish conformance requirements for exactly how a decoder would use the provided SEI message to produce the pictures it would display in response to receiving the SEI message. Instead, a core decoding process is applied as for ordinary picture decoding, and the GFV processing is applied as a post-process that is outside of the motion-compensated prediction loop. This approach offers decoder implementers the freedom to customize their implementations to suit their own needs or for specifications establishing such details to be developed separately in other organizations. To facilitate this approach, the draft syntax for the GFV SEI message supports the ability for an encoder to use Uniform Resource Identifiers (URIs) to identify such third-party specifications.
Although the proposed GFV SEI message can encapsulate various GFVC features into a unified syntax format, the existence of various GFVC algorithms can still result in mismatched encoder and decoder, where the encoder chooses to send a type of facial representation that is different from what the generator network at the decoder expects to use. In other words, various types of GFVC features with distinct characteristics may not be mutually compatible during the face generation process. To solve the issue of interoperability among different GFVC systems, JVET has included a feature translator network [24] into the usage model for the GFV SEI message. The translator network can convert received GFVC features into ones that are specific to the decoder’s preferred model for interpretation. As such, different GFVC features can be effectively adapted to one another while retaining promising coding performance.
In addition, limited by the capabilities of generative models and the internal GFVC mechanisms (e.g., unchanged textures provided by a base picture and sometimes insufficient motion representation due to compact representation), the reconstructed face video quality sometimes may not be robust. Such instability in reconstruction quality mainly includes objectionable background distortions, and/or inaccurate local motions in the mouth and/or eyes, which may damage the feasibility of practical applications. To improve reconstruction robustness, the JVET GFVC AhG has investigated a series of image processing techniques for signal fusion and enhancement, including a scalable representation and layered reconstruction scheme [35], background segmentation/inpainting/fusion operations [27, 25], and pupil position redirection [36].
III-C Encoding/Decoding Processes
The SEI-based GFVC workflow is shown in Fig. 1, including the encoding and decoding processes. The SEI-based GFVC encoder performs feature analysis on the input face video to obtain compact facial representations via the analysis model, and then encapsulates these learned compact parameters according to the syntax defined by the GFV SEI message. Afterwards, these GFV SEI messages are packed together with the coded base pictures into VVC coded bitstreams. To further conform to the VVC Main Profile, the base pictures are coded in the 4:2:0 format.
The decoder receives the bitstream to reconstruct the face video. In particular, a VVC decoder with support for the GFV SEI message is utilized to decode the base picture and parse the GFV SEI message. Further, the parsed GFV SEI parameters and the VVC-decoded base pictures are jointly fed into the synthesis model (i.e., deep generative network) to render the reconstructed video. It should be mentioned that a software implementation that streamlines all encoding and decoding processes has been proposed and made available by JVET GFVC AhG for further experimentation [37]. Decoders that do not support the GFV SEI message would simply decode and display the result of the ordinary standardized core hybrid decoding process.
IV Experiments
IV-A Experimental Settings
IV-A1 Implementation Details
The GFV SEI message has been included in the working draft of VSEI version 4 [12] and its implementation on top of the VVC reference software has been made available in the VTM SEI branch 111VTM Software with GFV SEI Implementation. In addition, as the output of the JVET GFVC AhG activities, a GFVC software tool has been developed and made available on a GFVC git branch 222JVET AhG16 GFVC Software Tools. Additionally, python scripts are provided to link the GFVC software tool with the GFV SEI message implementation, such that the entire GFVC encoding/decoding processes using the GFV SEI approach can be executed in a streamlined manner.
IV-A2 Testing Conditions
Experiments of GFVC using a GFV SEI message implementation were conducted under the JVET GFVC AhG test conditions [9]. Four representative GFVC algorithms (i.e., CFTE [4], FV2V [3], DAC [2] and FOMM [29]) and a lightweight GFVC version CFTE-Lite are selected as testing algorithms. All testing sequences 333JVET AhG16 Testing Sequences are available in two resolutions 256256 and 512512. Herein, these test sequences span considerable diversity of face video content in terms of age, gender, ethnicity, appearance, expression, head position, head motion, camera motion, background, etc. They are pre-processed (i.e., cropped) from open-source high-quality high-resolution video, and therefore have smaller resolutions relative to the original video and have squared aspect ratio. The detailed experimental settings are listed as follows:
For an anchor, VTM-22.2 reference software for VVC is used under low-delay B (LDB) configuration with quantization parameter (QP) set to 37, 42, 47 and 52.
For the tests, the first picture in each sequence is taken as the base picture and coded as an I picture using VTM-22.2 with QP values of 22, 32, 42 and 52. All other pictures are coded with the SEI-based GFVC approach, where the compact feature parameters are signaled via the GFV SEI message.
For objective quality evaluation of the reconstructed face video, we adopt Deep Image Structure and Texture Similarity (DISTS) [38] and Learned Perceptual Image Patch Similarity (LPIPS) [39], where a smaller score of these two measures indicates higher image quality.
| Resolution | Side | Measure | VVC | CFTE | FV2V | DAC | FOMM | CFTE-Lite |
| 256×256 | Encoder | Params (M) | \ | 14.12 | 72.13 | 14.15 | 14.21 | 1.62 |
| kMACs/pixel | \ | 15.10 | 120.28 | 16.20 | 19.49 | 1.84 | ||
| FPS | 0.36 | 16.93 | 14.12 | 16.89 | 15.08 | 16.45 | ||
| Decoder | Params (M) | \ | 44.11 | 53.32 | 45.80 | 45.80 | 6.03 | |
| kMACs/pixel | \ | 1057.71 | 3211.12 | 1039.71 | 1039.71 | 172.40 | ||
| FPS | 504.29 | 24.09 | 10.11 | 23.23 | 21.58 | 27.03 | ||
| 512×512 | Encoder | Params (M) | \ | 14.12 | 72.13 | 14.15 | 14.21 | \ |
| kMACs/pixel | \ | 3.77 | 120.28 | 4.05 | 4.87 | \ | ||
| FPS | 0.11 | 9.08 | 8.07 | 9.07 | 8.56 | \ | ||
| Decoder | Params (M) | \ | 44.11 | 53.32 | 45.80 | 45.80 | \ | |
| kMACs/pixel | \ | 389.13 | 927.48 | 384.63 | 384.63 | \ | ||
| FPS | 156.52 | 15.39 | 7.80 | 14.51 | 14.34 | \ |
IV-B Performance Comparisons
Fig. 2 shows that promising rate-distortion performance (i.e., Rate-DISTS and Rate-LPIPS) can be achieved with the proposed SEI-based GFVC approach compared to the traditional VVC standard in higher bitrate ranges for the 256256 resolution. It can be also seen that with the increase of face video resolution to 512512, the proposed SEI-based GFVC approach can achieve coding performance benefits for a wider range of bit rates. This is because the overhead of sending facial parameters via SEI messages becomes smaller relative to the VVC-coded base pictures as the video resolution increases. Such performance results confirm that SEI-based GFVC can provide promising coding performance for face video with resolutions at or above standard definition television resolution.
Fig. 3 provides subjective quality comparisons using VVC and the proposed SEI-based GFVC approaches at similar bit-rates. At ultra-low bitrates, VVC reconstructed face images are seriously distorted, whereas the proposed SEI-based approaches can still reconstruct vivid faces with higher fidelity and provide much better texture details.
IV-C Functionalities Examples
In addition to ultra-low bandwidth video communication, this SEI-based GFVC approach enables various purposes like user-specified animation/filtering and metaverse-related applications, with capabilities that cannot be directly supported by the core VVC codec. As illustrated in Fig. 4, the proposed approach gives the sender control over the quality of animated or filtered face pictures, and it can also support parameter modification in the GFV SEI message for expression or posture interactions in a virtual world.
IV-D Complexity Analysis
To show the model complexity and coding efficiency of SEI-based GFVC approaches, we compare the number of model parameters (Params/M), number of operations measured by Multiply–Accumulate (kMACs/pixel) as well as actual inference efficiency (FPS). Herein, we employ four original GFVC models (i.e., CFTE/FOMM/DAC/FV2V) and a lightweight GFVC model (i.e., CFTE-Lite) in this test. The overall encoding/decoding processes are both executed on Tesla V100-SXM2-32GB with 24 core CPUs (Intel(R) Xeon(R) Platinum 8163 CPU @ 2.50GHz). In addition, the testing sequence has 250 frames at the two different resolutions 256256 and 512512. To avoid the bias, the actual inference time at the encoder and decoder sides are both averaged with different QP settings according to JVET GFVC AhG test conditions [9].
As shown in Table II, compared with the traditional VVC codec, SEI-based GFVC schemes can significantly reduce the actual inference time of encoder side, while the computational complexity and inference efficiency are both increased for decoding. As for these GFVC models, their decoders are usually faced with more parameters and higher complexity than their encoders, since these decoders are designed with sophisticated architecture for accurate motion estimation and vivid signal generation. In addition, with the increase of video resolution, the value of kMACs-per-pixel can be reduced. Moreover, the lightweight GFVC version CFTE-Lite can achieve good trade-offs among model size, complexity and actual inference time whilst maintaining promising RD performance. In particular, CFTE-Lite can greatly reduce the computational complexity of CFTE encoder (from 15.10 kMACs/pixel to 1.84 kMACs/pixel) and decoder (from 1057.71 kMACs/pixel to 172.40 kMACs/pixel) at the video resolution of 256256.
IV-E Latency Analysis
Herein, we have provided latency evaluation results and analyzed the latency of different SEI-based GFVC approaches from the following aspects: average bit ratio (base picture v.s. all pictures) and bit fluctuation at picture level.
Table III provides the average bit ratio that is used for the base picture and all pictures at different base picture QP settings. It can be revealed that the bit consumption of SEI-based GFVC approaches is more susceptible to the QP and resolution of base picture instead of GFV SEI messages. Specifically, as the QP of the base picture decreases or the resolution of the base picture increases, the bit ratio of base picture v.s. all pictures will both gradually increase.
In addition, Fig. 5 shows picture-level bit fluctuation for SEI-based GFVC approaches for two test sequences with 250 frames but different resolutions. It should be noted that the bit consumption of base picture is not counted but the bit consumption of its SEI messages is counted. As illustrated in Fig. 5, the proposed SEI-based GFVC approaches can stably allocate coding bits to all pictures without any obvious bit fluctuation. As such, this proposed SEI-based GFVC approach can well support ultra-low bitrate communication.
| SEI-based GFVC Methods | Resolution | GFVC Bit Ratio (Base Picture v.s. All Pictures) | |||||||||
| CFTE | 256×256 | 51% | 30% | 15% | 8% | ||||||
| 512×512 | 72% | 48% | 25% | 12% | |||||||
| FV2V | 256×256 | 53% | 31% | 16% | 8% | ||||||
| 512×512 | 73% | 49% | 25% | 12% | |||||||
| DAC | 256×256 | 58% | 36% | 19% | 10% | ||||||
| 512×512 | 78% | 54% | 30% | 15% | |||||||
| FOMM | 256×256 | 68% | 45% | 24% | 12% | ||||||
| 512×512 | 73% | 49% | 27% | 13% | |||||||
| CFTE-Lite | 256×256 | 53% | 30% | 15% | 8% | ||||||
V Conclusions and Discussions
Generative AI has dramatically changed the acquisition of digital content and driven forward the innovative progress of visual data compression. However, these promising generative coding technologies face the challenge of providing a standardized way to represent various feature formats, thus greatly limiting their practical deployment and breadth of applications. To address this, this paper discusses how to define a unified high-level syntax structure for different GFVC representations and insert it into a VVC bitstream as SEI messages, and can incorporate different GFVC algorithms in a standardized workflow. The simulation results demonstrate the effectiveness of the proposed SEI-based GFVC approach in delivering promising performance and diverse functionalities.
As for potential future work related to GFVC technique and its standardization, more robust model and stable reconstruction quality may improve the feasibility of practical applications. Besides, it is necessary to further reduce model complexity especially for the GFVC decoder and reduce deployment cost especially on mobile devices. Another promising direction is to systematically investigate the quality evaluation of generative coding, since the existing pixel-level or perceptual-level quality measures are not entirely suitable to reflect the quality of reconstructed video. Furthermore, it is envisioned that the proposed approach could inspire the development of future generative video coding technologies beyond face video. In particular, it is straightforward to extend the proposed standardization scheme from talking face video to human moving body video or even some natural-scene video that possess strong prior knowledge in structure of motion and texture, which can facilitate the generative video coding towards diverse scenarios and collaborated communication.